14:00: system is up since the morning. We close the loop on LGSW only. We check QC behaviour because we have problems keeping the TT loop closed on that. One quadrant shows bad behaviour (less light than others).

We acquire QC dark: [187, 130, 141, 107], 40s integration. During May run we had [230, 110, 116, 110], that is incompatible.

LUCI rotator is at 300deg (equivalent to -60deg). In this way we have the XY stages directions parallel to the CCD47 pixels grid.

We acquire a Flat field using the dome light. We compute a relative efficiency of [83, 100, 87, 11]%. The APD#4 shows a very low throughput... Gilles goes up to check how the fibers are plugged. He notices that the fiber on APD#4 is not properly plugged (they have a key and it was not properly rotated to match the key into its slot).

Second Flat field measurement: [83, 98, 86, 100]% this is compatible with previous measurements.

16:30: we discover that fibers 2 and 1 are swapped on the AGW patch panel.

17:00: fibers adjusted, we close the TT loop on the QC

After the handover we try a Pockels Cells Test:

TN Sup
'20150630_025624','20150630_025749' [107, 111, 138]
'20150630_025951','20150630_030117'
-- MarcoBonaglia - 30 Jun 2015
Topic revision: r2 - 05 Aug 2015, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback